OpenStudy (bloomlocke367):

What is the smallest possible value of theta for which sin(2(theta)) = cos(4(theta))

OpenStudy (bloomlocke367):

@IrishBoy123 @nincompoop @pooja195

OpenStudy (bloomlocke367):

@aaronq

OpenStudy (irishboy123):

why not sub \(\phi = 2 \theta\) first, so you are into simple double angle formulae.....see if it makes sense and if there is an answer....

OpenStudy (bloomlocke367):

What? I don't understand... @IrishBoy123

OpenStudy (mathmale):

Bloom: "What is the smallest possible value of theta for which sin(2(theta)) = cos(4(theta))"\[\sin 2 \theta=\cos 4\theta\]

OpenStudy (mathmale):

can be solved in several ways. One way would be to graph both functions on the same set of axes and determine the first angle theta at which the two graphs intersect. There is a "double angle formula" for the cosine function. supposing we wanted cos 2x, we could write that as cos (x+x). cos (x+x) = cos x cos x - sin x sin x, or \[\cos^2x-\sin^2x.\]

OpenStudy (mathmale):

This, in turn, can be re-written as \[(1-\sin^2 x)-\sin^2x, .or.1-2\sin^2x. \]

OpenStudy (mathmale):

This is a variation of what IrishBoy was asking you to do.

OpenStudy (mathmale):

See whether you can prove to yourself that \[\cos4 \theta=\cos(2\theta+2\theta)=1-\sin^2 2\theta\]

OpenStudy (mathmale):

Now let's go back to the original problem: \[\sin 2\theta=\cos 4 \theta\]

OpenStudy (mathmale):

and replace the right side with its equivalent\[1-\sin^22\theta\]

OpenStudy (mathmale):

Then you have sin2theta on the left equal to 1-(sin 2theta)^2.\[\sin2\theta=1-\sin^22\theta\]

OpenStudy (mathmale):

Write this in the form of a quadratic equation. (Move the right 2 terms to the left side and replace them with zero on the right side.) Can you now solve this quadratic equation?

OpenStudy (irishboy123):

Hi Bloom! i mean, if you re-write it as: \( \sin \phi = cos2 \phi\), where \(2 \theta = \phi\), then you can then say stuff like: \( \sin \phi = 1 - 2 \sin^2 \phi\) so you can explore it as a quadratic, as @mathmale suggests.......

OpenStudy (bloomlocke367):

Well, the issue is, we can't use a calculator .-. it's for a math competion

OpenStudy (bloomlocke367):

I'm looking at old problems so yea

OpenStudy (irishboy123):

now we dig some more from \(\sin \phi = 1 - 2 \sin^2 \phi\) we say that: \(a = 1 - 2 a^2\) can you solve that for a??

OpenStudy (bloomlocke367):

one moment.

OpenStudy (irishboy123):

all we've done is say that \(a = \sin \phi = \sin 2 \theta\)

OpenStudy (bloomlocke367):

-1, and 1/2

OpenStudy (irishboy123):

think that's good so that's it solved for "a" but we've said \(a = \sin \phi = \sin 2 \theta\) so \(a = \sin \phi = \sin 2 \theta = -1\) and \(a = \sin \phi = \sin 2 \theta = 1/2\)

OpenStudy (irishboy123):

you can solve that for \(\phi\), then \(\theta\).... .... or for \(2 \theta\),... and you're done.

OpenStudy (mathmale):

There's certainly more than one way in which to skin a cat. However, I'd still advocate using my model, which has resulted in a quadratic equation. \[\sin2\theta=1-\sin^22\theta\]

OpenStudy (mathmale):

Moving both terms from the right side to the left side,\[\sin^22 \theta+\sin2\theta-1=0\]

OpenStudy (mathmale):

Letting\[\sin 2\theta = x\]

OpenStudy (mathmale):

this boils down to x^2+x-1=0. Determine whether or not this has any real roots. If c is a real root, then solve the equation c=sin 2theta. Remember, we want the smallest theta will satisfy the original equation.